OpenStudy (anonymous):

Find 8th term in GEOMETRIC sequence... please help?

OpenStudy (anonymous):

...Confused as to how to start this?

OpenStudy (anonymous):

looks like you are multiplying by 5 each time right?

OpenStudy (anonymous):

I think so @satellite73

OpenStudy (anonymous):

\[\large 4\times 5,4\times 5^2,4\times 5^3, \color{red}{4\times 5^4}\]

OpenStudy (anonymous):

2500 will be 8th term, right?

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (anonymous):

Thank you @satellite73 !

OpenStudy (anonymous):

yw